OpenStudy (anonymous):

state the conjugate of 6+sqrt2 and then multiply them together

OpenStudy (anonymous):

@jdoe0001 can you help me with this as well?

OpenStudy (jdoe0001):

conjugate of a pair, is the same exact pair, with a different sign in the middle so \(\bf a + b\), it's conjugate is \(\bf a - b\) \(\bf something - other\) conjugate is \(\bf something + other\) and so on, pretty much the same thing, if it has a + first, the conjugate will use the - if it has the - first, the conjugate will use the +

OpenStudy (anonymous):

so its 6-sqrt2?

OpenStudy (jdoe0001):

yes \(\bf 6+\sqrt{2}\qquad conjugate \implies 6-\sqrt{2}\\ \quad \\ \textit{now, conjugates multiplication ends up like}\qquad (a-b)(a+b) = a^2-b^2\\ \quad \\ thus\\ \quad \\ (6+\sqrt{2})(6-\sqrt{2})\implies 6^2-(\sqrt{2})^2\implies 6^2-\sqrt{2^2}\)

OpenStudy (anonymous):

okay, okay! Thanks, thats pretty easy

OpenStudy (jdoe0001):

yw